# Investigating potential causes of Sepsis with Bayesian network structure learning

### v1

@@ 6 Conclusion
The algorithms that perform structure learning can capture interrelations among variables. Since the interrelations captured are sensitive to the algorithms used, model averaging was performed to investigate how the overall graph learnt from data compares to clinical expertise. We find that the graphs learnt from data alone differ considerably from those produced through clinical expertise, but this does not necessarily imply that some of these graphs must be wrong or not useful. In other words, many of the edges produced by the algorithms are deemed clinically sensible, despite not being present in the knowledge graph we initially produced. This highlights the need to consider both expertise and data-driven approaches to derive new understanding.
@@
This study comes with some limitations that guide future research directions. Firstly, the data available for this analysis is a snapshot of hospital episodes. The clinical expert stated that it is possible that a patient with a history of Sepsis might be at a higher risk of re-occurrence of Sepsis. He quantified this belief with 40% confidence. A more suitable dataset and the application of structure algorithms capable of discovering structures from time series data could address this issue. Secondly, it is important to reiterate that the results presented in this study assume that the edges produced by the algorithms and the clinical expert represent causal relationships, which is needed to perform causal inference. Any conclusions drawn from this study should be mindful of this fact. Lastly, since the data used in this study comes from hospitals in England, future research could explore whether these results are consistent with those obtained from hospitals in other countries, potentially identifying important differences between countries exercising similar or different relevant practices, and enhancing the generalizability of the model.

### v2

@@ 6 Conclusions
The algorithms that perform structure learning can capture interrelations among variables. Since the interrelations captured are sensitive to the algorithms used, model averaging was performed to investigate how the overall graph learnt from data compares to clinical expertise. We find that the graphs learnt from data alone differ considerably from those produced through clinical expertise, but this does not necessarily imply that some of these graphs must be wrong or not useful. In other words, many of the edges produced by the algorithms are deemed clinically sensible despite not being present in the knowledge graph we initially produced. This highlights the need to consider both expertise and data-driven approaches to derive new understanding.
@@
This study comes with some limitations that guide future research directions. Firstly, the data available for this analysis is a snapshot of hospital episodes. The clinical expert stated that it is possible that a patient with a history of Sepsis might be at a higher risk of re-occurrence of Sepsis and quantified this belief with 40% confidence. Furthermore, the data used lacks relevant clinical features such as biomarkers, lab results or medication dosages, which could improve model performance. A richer dataset could enable us to incorporate levels of severity, past diagnosis information, or build time-varying features such as changes in lactate levels over time.
Secondly, it is important to reiterate that the results presented in this study assume that the edges produced by the algorithms and the clinical expert represent causal relationships, which is a necessary assumption for causal inference. Enforcing certain relationships based on clinical expertise to refine the structure learnt by the algorithms may be subject to other forms of bias, and the inclusion of a literature review aimed to mitigate this risk. Future work could explore more sophisticated ways to integrate expert knowledge, by expanding the datasets with richer clinical features, which would allow for more detailed causal representations, thereby improving the overall model performance.
Lastly, the missing data (0.6%) was handled using the well-established missForest R package. While the small percentage of missing values and the large sample size (1 million hospital episodes) suggest minimal impact on the study’s findings, alternative imputation methods were not explored. Moreover, since the data used in this study comes from hospitals in England, future research could explore whether these results are consistent with those obtained from hospitals in other countries, potentially identifying important differences between countries exercising similar or different relevant practices, and enhancing the generalisability of the model.